OpenStudy (anonymous):

PLEASEEEE. Solve the equation 2cosx - 1 = 0 for all values of x in the interval 0 degrees less than or equal to x less than 360 degrees

OpenStudy (anonymous):

cosx = 1/2 => x = 60 + 2K * 180 Thus x = 60

OpenStudy (anonymous):

thank you so much